When mineral deposits formed a coating 1mm thick on the inside of a pipe, the area through which fluid can flow was reduced by 20%. Find the original inside diameter of the pipe
Answer by edjones(8007)

pi*d=C
Let d=original inside diameter of the pipe
.8C=pi(d-.002)
.
=>C=pi*d
.8C=pi*d-.002*pi
--------------------------subtract
.2C=.002*pi
.
C=.01*pi
pi*d=.01*pi
d=.01 m or 1 cm
.
Ed